the area of a rectangle is 450 square inches. If the length is 24 inches, what is the width?
Found 2 solutions by texttutoring, fractalier:
Answer by texttutoring(324)

Area = length * width
width = Area / length
width = 450 / 24
width = 18.75

Answer by fractalier(6550)

Area = length x width
450 = 24 x width
width = 450 / 24 = 18.75